# LLM-Assisted Relevance Assessments: When Should We Ask LLMs for Help?

### v2

@@ Abstract.
Test collections are information retrieval tools that allow researchers to quickly and easily evaluate ranking algorithms. While test collections have become an integral part of IR research, the process of data creation involves significant effort in manual annotations, which often makes it very expensive and time-consuming. Thus, test collections could become too small when the budget is limited, which may lead to unstable evaluations. As a cheaper alternative, recent studies have proposed the use of large language models (LLMs) to completely replace human assessors. However, while LLMs may seem to somewhat correlate with human judgments, their predictions are not perfect and often show bias. Thus a complete replacement with LLMs is argued to be too risky and not fully reliable.
Thus, in this paper, we propose LLM-Assisted Relevance Assessments (LARA), an effective method to balance manual annotations with LLM annotations, which helps to build a rich and reliable test collection even under a low budget. We use the LLM’s predicted relevance probabilities to select the most profitable documents to manually annotate under a budget constraint. With theoretical reasoning, LARA effectively guides the human annotation process by actively learning to calibrate the LLM’s predicted relevance probabilities. Then, using the calibration model learned from the limited manual annotations, LARA debiases the LLM predictions to annotate the remaining non-assessed data. Empirical evaluations on TREC-7 Ad Hoc, TREC-8 Ad Hoc, TREC Robust 2004, and TREC-COVID datasets show that LARA outperforms alternative solutions under almost any budget constraint.
offline evaluation; relevance judgments; test collections
ccs: Information systems Test collections
@@ 1. Introduction
To address the dilemma between LLMs and humans, our work proposes LLM-Assisted Relevance Assessments (LARA) to achieve a reliable, yet budget-friendly annotation procedure. Given a limited budget, LARA uses LLMs to choose the most effective documents for manual annotation. Specifically, LARA uses LLMs’ output token probabilities to find their most uncertain predictions that should be manually annotated. Then, the rest of the documents (easy and more confident ones) can be judged by LLMs. While directly using LLM relevance probability as uncertainty is an effective method to find documents for manual annotation, we argue and show that these LLM predictions are often biased and are not optimal. Therefore, we propose an algorithm that actively learns to calibrate the LLM relevance probabilities, while continuously identifying the most effective documents to annotate manually. The remaining documents are then annotated using the calibrated LLM predictions, adjusted by the calibration model learned from human assessments. Thus, in our algorithm, calibrated LLM predictions guide manual annotations, and the gathered manual annotations in turn refine LLM predictions. Moreover, while most existing algorithms for efficient manual annotations are only suited for binary relevance labels (Cormack et al., 1998; Aslam et al., 2006; Li and Kanoulas, 2017), our LARA algorithm can deal with the general graded relevance framework (e.g., relevance levels 0/1/2), where the binary relevance setting is only a special case of our framework. Finally, we conduct comprehensive experiments on TREC-7 Ad Hoc, TREC-8 Ad Hoc, TREC Robust 2004, and TREC COVID datasets, where we show that our LARA algorithm almost constantly outperforms the tested alternative methods under various budget constraints. We are also the first to study and compare the performance of LLM-based methods and manual methods under different budget constraints.
@@ 2.1. Efficient Human Annotations
When a document collection is large, annotating all topic-document pairs available is difficult due to the immense human labor required. To solve this problem, various methods have been proposed to select subsets of documents for human annotation (Cormack et al., 1998; Aslam et al., 2006; Aslam and Yilmaz, 2007; Li and Kanoulas, 2017; Voorhees, 2018).
@@ 2.2. Automated Relevance Assessments
These studies have shown the capability of LLMs in relevance assessments and have found that LLMs somewhat correlate with the ground truth labels and often perform better than crowd workers (Thomas et al., 2024). However, the relevance predictions by LLMs are often biased. For example, Alaofi et al. (Alaofi et al., 2024) report that LLMs tend to over-generate relevance labels compared to the ground truth. If we were to build a test collection only using LLMs, we would not be able to avoid such bias, making a full replacement risky (Soboroff, 2024; Clarke and Dietz, 2024). Even if a well-performing prompt is found on a certain dataset, there is no guarantee that it will work on an unseen task or data that we want to use for a test collection. To minimize such risks posed by LLMs, our LARA algorithm effectively uses partial ground truth judgments collected under a limited budget to actively calibrate the LLM predictions, building a more reliable dataset.
@@ 3.2. A Naive Uncertainty-Based Approach
where $p_{\text{LLM}}(k|x)$ is the probability of the LLM assigning the relevance level $k$ to the question “How relevant is the document to the topic?”. We express the LLM probability set for data point $x_{i}$ as $\pi_{i}=\{\pi_{i}^{0},\pi_{i}^{1},...,\pi_{i}^{l}\}$. The prompts are provided in Appendix A.
Given these probabilities, one naive approach is to identify the data points on which LLM is the most uncertain and annotate those manually, which we call the Naive method. For a graded relevance setting, a simple heuristic is to first determine the most likely predicted level $k^{\prime}$ and the second most likely predicted level $s^{\prime}$ for each data point:
@@ 3.3. LLM-Assisted Relevance Assessments
Theorem 1 (Optimal Inspection Point).
Let $p(y=j|\pi)$ denote the true probability of relevance level $j$ given $\pi$. Let $k=\operatorname*{arg\,max}_{j}p(y=j|\pi)$ be the level with the highest probability and $s=\operatorname*{arg\,max}_{j\neq k}p(y=j|\pi)$ be the level with the second highest probability. Then, the greatest potential reduction in classification error is achieved by annotating the data point with the smallest margin
@@
where we suppose that the remaining data are classified into $\hat{y}=\text{argmax}_{j}p(y=j|\pi)$. We share the proof upon acceptance. Appendix B proves this via the maximum uncertainty principle.
@@ 4.2. Comparison Methods
We compare the LARA algorithm with several baseline approaches. We first introduce methods from previous studies that only use manual assessments. Note that since most of these existing manual methods are only suited for binary relevance levels, we binarize the relevance levels in the learning procedure by setting a threshold at $\frac{l}{2}$, but we keep the graded relevance levels for final annotation.
@@
The base prompt used in the experiments is provided in Appendix A. To study the effect of the prompt on LARA, we also share the results with the rationale prompt based on Upadhyay et al. (Upadhyay et al., 2024b) and the utility prompt based on Thomas et al. (Thomas et al., 2024) as an extensive study. The prompts are slightly modified from the original to better align with our task [^1] We share the full rationale and utility prompts upon acceptance.. Methods that use these prompts are named in the format MethodName(PromptType), like LARA(rationale) and LARA(utility), where the LARA methods here use LARA($n=N$) for the number of annotators.
@@ 5.1. Main Results
How does the number of annotators affect the performance of LARA? The LARA method with different numbers of annotators (i.e., LARA($n=1$), LARA($n=3$), and LARA($n=N$)), perhaps surprisingly, have almost identical performance on all datasets and budgets, with LARA($n=N$) slightly outperforming the other two in many scenarios. We conjecture that this occurs because as the number of annotators $n$ increases, the number of ground truth annotations would be distributed fairly among each group of topics. This allows each topic to have similar qualities of evaluation results, which helped increase the overall performance. In contrast, when the number of annotators $n$ is small, although the algorithm would be able to choose the most likely-uncertain data point from a wider range of topics, the evaluation performance could vary greatly from topic to topic, since some topics could have very accurate evaluations while others could have poor performance due to differences in manual annotation distribution. This interesting result supports the practicality of our LARA method. We would generally recommend using LARA($n=N$) since it can handle any number of assessors.
@@
How does the performance of the LLM affect the quality of LARA? We also show the results of LLM-based methods using Llama-3.1-8B-Instruct, a lightweight variant of Llama-3.1-70B-Instruct. Comparing the LLM-Only methods, we first know that the 8B model performs constantly worse than the 70B model. Nonetheless, while LARA (Llama-8B) generally never outperforms LARA with the 70B models, LARA (Llama-8B) outperforms the rest of the 8B-based models, again showing the effectiveness of the algorithm. Compared to the existing human-based methods, LARA (Llama-8B) performs better under low-budget constraints for all datasets. Thus, when building a dataset under a low budget, LARA is a good option even with a small LLM, but we definitely recommend using a larger and better LLM if available.
@@ 6. Discussions and Future Directions
However, it may be beneficial to fine-tune the LLM pre-deployment, using historical or domain-specific data. Such pre-deployment tuning could yield a more accurate starting point for LARA, as we see from the results that the quality of LLMs affects the performance of LARA. Exploring the trade-offs between continuous retraining, pre-deployment fine-tuning, and on-the-fly calibration is an interesting avenue for future research. As future work, it would also be valuable to investigate a LARA-like hybrid method on other annotation tasks like nugget matching (Takehi et al., 2023; Nenkova and Passonneau, 2004), which aims to annotate whether some information nugget can be obtained from a text, and e-Discovery (Oard et al., 2018; Cormack and Grossman, 2014), which aims to accurately collect relevant documents for use in legal cases.
@@ 7. Conclusions
This paper investigates an effective and efficient way of building a test collection, particularly using both LLM and limited manual assessments. Relying only on limited manual assessments leads to unstable evaluations due to limited collection size, and relying only on LLM judgments is risky and unreliable. Thus, we propose LLM-Assisted Relevance Assessments (LARA) to build a risk-minimized yet large test collection. LARA actively learns to calibrate LLM relevance predictions by strategically selecting documents for human annotation. This learned calibration model guides which documents should be manually reviewed and corrects the LLM predictions for the remaining documents. Experiments on various datasets demonstrate that LARA consistently outperforms all the tested alternatives under almost any budget constraint, suggesting that LARA could help venues like TREC, CLEF, NTCIR, and INEX obtain large, stable, and reliable test collections even when the budget for manual annotations is limited.
Appendix A Prompt
We used the following base prompt for graded relevance, where {description} and {narrative} are given by the NIST assessors.
LLM Prompt Consider the following web page content:
-BEGIN WEB PAGE CONTENT-
{document} -END WEB PAGE CONTENT- A person has typed “{topic}” into a search engine. This person’s intent of this query was “{description}”. Instruction:
Answer if the web content is relevant to the query.
{narrative}. Answer {2} for highly relevant, {1} for somewhat or partially relevant, or {0} for not relevant. Your answer:{
For binary relevance assessments, the final two lines are:
Answer yes or no. Your answer:
Appendix B Proof of Theorem 1
Proof.
Let $p(y=j|\pi)$ be the true relevance probability of relevance level $j$. The Bayes-optimal classifier would predict :
$k=\operatorname*{arg\,max}_{j}p(y=j|\pi).$
Its error is:
$\text{Error}(\pi)=1-p(y=k|\pi).$
Define:
$s=\operatorname*{arg\,max}_{j\neq k}p(y=j|\pi),\quad m=p(y=k|\pi)-p(y=s|\pi)\geq 0.$
Maximizing $\text{Error}(\pi)$ is equivalent to minimizing $p(y=k|\pi)$ under the constraint $p(y=k|\pi)\geq p(y=s|\pi)$. This minimum is achieved at $m=0$, i.e.,
$p(y=k|\pi)=p(y=s|\pi).$
At this point, the classifier is most uncertain. Thus, minimizing $m$ yields the greatest potential reduction in classification error. ∎

### v4

@@ Abstract.
Test collections are information retrieval tools that allow researchers to quickly and easily evaluate ranking algorithms. While test collections have become an integral part of IR research, the process of data creation involves significant manual annotation effort, which often makes it very expensive and time-consuming. Consequently, test collections could become too small when the budget is limited, which may lead to unstable evaluations. As a cheaper alternative, recent studies have proposed the use of large language models (LLMs) to completely replace human assessors. However, while LLMs seem to somewhat correlate with human judgments, their predictions are not perfect and often show bias. Thus, a complete replacement with LLMs is argued to be too risky and not fully reliable.
In this paper, we propose LLM-Assisted Relevance Assessments (LARA) [^1] Our code is available at https://github.com/RikiyaT/LARA, an effective method to balance manual annotations with LLM annotations, which helps to build a rich and reliable test collection even under a low budget. We use the LLM’s predicted relevance probabilities to select the most profitable documents to manually annotate under a budget constraint. With theoretical reasoning, LARA effectively guides the human annotation process by actively learning to calibrate the LLM’s predicted relevance probabilities. Then, using the calibration model learned from the limited manual annotations, LARA debiases the LLM predictions to annotate the remaining non-assessed data. Experiments on TREC-7 Ad Hoc, TREC-8 Ad Hoc, TREC Robust 2004, and TREC-COVID datasets show that LARA outperforms alternative solutions under almost any budget constraint. While the community debates humans vs. LLMs in relevance assessments, we contend that, given the same amount of human effort, it is reasonable to leverage LLMs.
offline evaluation; relevance judgments; test collections; large language models; active learning
journalyear: 2025copyright: licensedusgovmixedconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3729916isbn: 979-8-4007-1592-1/2025/07ccs: Information systems Test collectionsccs: Information systems Relevance assessmentccs: Computing methodologies Active learning
@@ 1. Introduction
To address the dilemma between LLMs and humans, our work proposes LLM-Assisted Relevance Assessments (LARA) to achieve a reliable, yet budget-friendly annotation procedure. Given a limited budget, LARA uses LLMs to choose the most effective documents for manual annotation. Specifically, LARA uses LLMs’ output token probabilities to find their most uncertain predictions that should be manually annotated. Then, the rest of the documents (easy and more confident ones) can be judged by LLMs. While directly using LLM relevance probability as uncertainty is an effective method to find documents for manual annotation, we argue and show that these LLM predictions are often biased and are not optimal. Therefore, we propose an algorithm that actively learns to calibrate the LLM relevance probabilities, while continuously identifying the most effective documents to annotate manually. The remaining documents are then annotated using the calibrated LLM predictions, adjusted by the calibration model learned from human assessments. Thus, in our algorithm, calibrated LLM predictions guide manual annotations, and the gathered manual annotations in turn refine LLM predictions. Moreover, while most existing algorithms for efficient manual annotations are only suited for binary relevance labels (Cormack et al., 1998; Aslam et al., 2006; Li and Kanoulas, 2017), LARA can deal with the general graded relevance framework (e.g., relevance levels 0/1/2), where the binary relevance setting is only a special case of our framework. Finally, we conduct comprehensive experiments on TREC-7 Ad Hoc, TREC-8 Ad Hoc, TREC Robust 2004, and TREC COVID datasets, where we show that our LARA algorithm almost constantly outperforms the tested alternative methods under various budget constraints. We are also the first to study the performance of LLM-based methods and manual methods under different budget constraints.
@@ 2.1. Efficient Human Annotations
When a document collection is large, annotating all topic-document pairs available is difficult due to the immense human labor required. To solve this problem, various methods have been proposed to select subsets of documents for human annotation (Cormack et al., 1998; Aslam et al., 2006; Aslam and Yilmaz, 2007; Li and Kanoulas, 2017; Voorhees, 2018; Ganguly and Yilmaz, 2023).
@@ 2.2. Automated Relevance Assessments
These studies have shown the capability of LLMs in relevance assessments and have found that LLM judgements somewhat correlate with the ground truth labels and often perform better than crowd workers (Thomas et al., 2024). However, the relevance predictions by LLMs are often biased. For example, Alaofi et al. (Alaofi et al., 2024) report that LLMs tend to over-generate relevance labels compared to the ground truth. If we were to build a test collection only using LLMs, we would not be able to avoid such bias, making a full replacement risky (Soboroff, 2024; Clarke and Dietz, 2024). Even if a well-performing prompt is found on a certain dataset, there is no guarantee that it will work on an unseen task or data that we want to use for a test collection. To minimize such risks posed by LLMs, our LARA algorithm effectively uses partial ground truth judgments collected under a limited budget to actively calibrate the LLM predictions, building a more reliable dataset. Fine-tuned LLMs have shown improvement in accuracy (Meng et al., 2024), but have crucial limitations in annotation latency, as we will discuss in Section 6.
@@ 3. LLM-Assisted Relevance Assessments
[…]
Four scatter–histogram panels comparing LLM-predicted relevance probabilities with true relevance on four TREC test collections. In every panel, grey bars show the distribution of LLM outputs, a blue line plots the empirical true-relevance rate, and a dotted 45-degree line marks perfect calibration; the blue line falls well below the dotted line except near probability 1.0, indicating the LLM is generally over-confident.
@@ 3.2. A Naive Uncertainty-Based Approach
where $p_{\text{LLM}}(k|x)$ is the probability of the LLM assigning the relevance level $k$ to a document given a query. We express the LLM probability set for data point $x_{i}$ as $\pi_{i}=\{\pi_{i}^{0},\pi_{i}^{1},...,\pi_{i}^{l}\}$.
Given these probabilities, one naive approach is to identify the data points on which LLM is the most uncertain and annotate those manually, which we call the Naive method. A simple heuristic is to first determine the most likely predicted level $k^{\prime}$ and the second most likely predicted level $s^{\prime}$ for each data point
@@ 3.3. LLM-Assisted Relevance Assessments
Theorem 1 (Optimal Inspection Point [^2] Proof can be found at https://github.com/RikiyaT/LARA/blob/main/docs/proof.md).
Let $p(y=j|\pi)$ denote the probability of the true relevance level being $j$ given LLM probabilities $\pi$. Let $k=\operatorname*{arg\,max}_{j}p(y=j|\pi)$ be the level with the highest probability and $s=\operatorname*{arg\,max}_{j\neq k}p(y=j|\pi)$ be the level with the second highest probability. Then, the greatest potential reduction in classification error is achieved by annotating the data point with the smallest margin
@@
where the remaining data are classified into $\hat{y}=\text{argmax}_{j}p(y=j|\pi)$.
@@ 4.2. Comparison Methods
We compare the LARA algorithm with several baseline approaches. We first introduce methods from previous studies that only use manual assessments. Note that since most of these existing manual methods are only suited for binary relevance levels, we binarize the relevance levels in the learning procedure by setting a threshold at $\frac{l}{2}$, but we keep the graded relevance levels in their final annotations.
@@
To study the effect of the prompt on LARA, we test three different prompts [^3] Prompts are available at https://github.com/RikiyaT/LARA#-prompt-templates. In addition to our simplest base prompt, we also share the results with the rationale prompt based on Upadhyay et al. (Upadhyay et al., 2024b) and the utility prompt based on Thomas et al. (Thomas et al., 2024) as an extensive study. The prompts are slightly modified from the original to better align with our task. Methods that use these prompts are named in the format MethodName(PromptType), like LARA(rationale) and LARA(utility), where the LARA methods here use LARA($n=N$) for the number of annotators.
@@
[…]
Four line charts showing overlap score versus manual‐annotation budget ratio for three methods (Naive, Random, LARA) on TREC-7, TREC-8, Robust04 and TREC-COVID. In every dataset the blue LARA curve stays consistently above the green Naive and orange Random curves, and the performance gap widens as budget increases, indicating LARA yields the fewest LLM-annotation errors.
@@ 5.1. Main Results
How does the number of annotators affect the performance of LARA? The LARA method with different numbers of annotators (i.e., LARA($n=1$), LARA($n=3$), and LARA($n=N$)), perhaps surprisingly, have almost identical performance on all datasets and budgets, with LARA($n=N$) slightly outperforming the other two in many scenarios. We conjecture that this occurs because as the number of annotators $n$ increases, the number of ground truth annotations would be distributed fairly among each group of topics. This allows each topic to have similar qualities of evaluation results, which helped increase the overall performance. In contrast, when the number of annotators $n$ is small, although the algorithm would be able to choose the most likely-uncertain data point from a wider range of topics, the evaluation performance could vary greatly from topic to topic, since some topics could have very accurate evaluations while others could have poor performance due to differences in manual annotation distribution. This interesting result supports the practicality of our LARA method. We would thus recommend using LARA($n=N$) since it can handle any number of assessors.
@@
How does the performance of the LLM affect the quality of LARA? We show the results of LLM-based methods using Llama-3.1-8B-Instruct, a lightweight variant of Llama-3.1-70B-Instruct. Comparing the LLM-Only methods, we first know that the 8B model performs constantly worse than the 70B model. Nonetheless, while LARA (Llama-8B) generally never outperforms LARA with the 70B model, LARA (Llama-8B) outperforms the rest of the 8B-based methods, again showing the effectiveness of LARA. Compared to the existing manual methods, LARA (Llama-8B) performs better under low-budget constraints for all datasets. Thus, when building a test collection under a low budget, LARA is a good option even with a small LLM, but we recommend using a better LLM if available.
@@ 6. Discussions and Future Directions
Nonetheless, it may be beneficial to fine-tune the LLM pre-deployment, using historical or domain-specific data (Meng et al., 2024). Such pre-deployment tuning could yield a more accurate starting point for LARA, as we see from the results that the quality of LLMs affects the performance of LARA. Exploring the trade-offs between continuous retraining, pre-deployment fine-tuning, and on-the-fly calibration is an interesting avenue for future research. As future work, it would also be valuable to investigate a LARA-like hybrid method on other annotation tasks like nugget matching (Takehi et al., 2023; Nenkova and Passonneau, 2004), which aims to annotate whether some information nugget can be obtained from a text, and e-Discovery (Oard et al., 2018; Cormack and Grossman, 2014), which aims to accurately collect relevant documents for use in legal cases.
@@ 7. Conclusions
This paper investigates an effective and efficient way of building a test collection, particularly using both LLM and limited manual assessments. Relying only on limited manual assessments leads to unstable evaluations due to limited collection size, and relying only on LLM judgments is risky and unreliable. Thus, we propose LLM-Assisted Relevance Assessments (LARA) to build a risk-minimized yet large test collection. LARA actively learns to calibrate LLM relevance predictions by strategically selecting documents for human annotation. This learned calibration model guides which documents should be manually reviewed and corrects the LLM predictions for the remaining documents. Experiments on various datasets demonstrate that LARA consistently outperforms all the tested alternatives under almost any budget constraint, suggesting that LARA could help venues like TREC, CLEF, and NTCIR obtain large, stable, and reliable test collections even when the budget for manual annotations is limited.